OpenStudy (anonymous):

Factor completely 2x^2+7x+6

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i am solving hold on plz

OpenStudy (anonymous):

No, I meant what did you say sorry for?

OpenStudy (anonymous):

i meant to put that in chat lol

OpenStudy (anonymous):

2x^2+7x+6 => find two numbers that multiply to 12, add to 7: 3 & 4? = 2x^2 + 4x + 3x + 6 = 2x(x + 2) + 3(x + 2) = (x + 2)(2x + 3)